# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 5.5 Testing and Debugging
Jupyter notebooks are widely used for prototyping, offering rapid iteration and immediate feedback [34]. However, testing in notebooks differs from traditional software testing, as it may require verifying code correctness at the cell level rather than treating the entire notebook as a single executable unit. Due to the lack of standardized testing practices, data scientists have adopted various approaches. For example, some embed test cases within the notebook, while others use separate test notebooks to verify the correctness [21]. This section describes studies related to the testing of notebooks, including empirical studies on bug analysis and bug detection.
@@ 5.5.1 Empirical Studies on Testing Notebooks
De Santana et al. [26] empirically analyzed the problems and challenges of identifying, diagnosing and resolving bugs in Jupyter notebooks. They presented the first comprehensive study of bugs in Jupyter Notebook projects, analyzing bugs in GitHub repositories and Stack Overflow posts, and interviews with data scientists. The study identifies eight types of bugs that are found in notebooks along with their studies. Shome et al. [147] investigated feedback mechanisms within ML-focused Jupyter notebooks and presented a taxonomy of explicit (assertions) and implicit (print statements and last cell outputs) feedback approaches. Their study also provided practical recommendations to encourage explicit testing for minimizing technical debt and improving reproducibility. Wang et al. [182] analyzed crashes in ML notebooks and categorized predominant exceptions such as NameError and ValueError. In addition, they reported that some root causes of these exceptions are related to the features of Jupyter Notebook, such as out-of-order cell executions and errors propagating from previous cells.
@@
Processing bugs Issues in memory or computation, i.e., memory overflow, data loss, or performance degradation. [155, 103, 197, 181, 196]
Implementation bugs General coding errors that can cause incorrect results or runtime errors, i.e., syntax, logic, variables, or algorithms-related bugs. [109, 12, 47, 45]
Cell defect Bugs related to notebook cell renderings, i.e., code cells, Markdown, or output interactions. [34]
Environments and settings bugs Problems from misconfigured environments or dependencies that interrupt notebook execution, i.e., missing libraries or lack of dependencies. [12]
@@ 5.5.2 Detecting Bugs
Detecting bugs or errors in Jupyter Notebook presents unique challenges due to nonlinear workflows and hidden states, distinguishing them from traditional programming environments [132]. Xin et al. [196] developed a framework to detect anomalies and identify their root causes by combining provenance data with performance metrics. Since Jupyter notebooks do not offer a built-in debugger, the study also presents strategies for debugging Jupyter notebooks by identifying cell workflows and the root cause of errors. Another bug-detecting study introduced a tool called Vizier, which helps debug notebooks by maintaining a complete version history of notebooks, cells, and datasets, while tracking potential errors through fine-grained provenance [12]. This tool maintains a complete version history for each notebook and allows reproducible data manipulation through a spreadsheet mode [12].
@@
Another method to detect potential bugs in notebooks is to focus on inconsistencies in variable assignments, output visualization, or logical operations within the notebook [109]. For example, output inconsistency occurs when the output of a notebook cell does not match the expected output. An implementation of such an inconsistency-based approach is the nbval notebook validation plugin [34]. nbval detects bugs in Jupyter notebooks by automatically executing each code cell and comparing its output with previously stored results. If there is any deviation, such as an unexpected change in numerical values, formatting differences, or execution failures, nbval reports a test failure. Wang et al. [181] showed that traditional static analyzers frequently fail to detect runtime tensor shape mismatch bugs in ML codes (particularly TensorFlow) within notebooks. Their study demonstrates how incorporating runtime information improves static analysis capabilities for detecting such common ML-related notebook bugs.
@@
Another type of inconsistency bug is the name-value inconsistency, which occurs when the name of a variable does not accurately reflect its assigned value. Unlike traditional software development environments, notebooks allow for nonlinear execution and frequent variable reuse, increasing the risk of such name-value inconsistencies [109]. For instance, naming a variable log_file while storing a list of all files in a directory can cause a name-value inconsistency bug, as a developer might assume that log_file stores a single file name, but it actually contains a list. To avoid confusion, the variable could be renamed to log_files or log_file_list. Patra and Pradel [109] developed Nalin, an automated tool for detecting name-value inconsistencies. This tool uses a neural model to predict whether a variable’s name and its assigned value are consistent. Nalin employs dynamic program analysis alongside deep learning to monitor variable activities throughout the execution process.
@@
Recently, advancements in LLMs have created promising opportunities for automated error detection and debugging. Grotov et al. [47] performed a detailed analysis of prevalent notebook error patterns and proposed an iterative, LLM-based method for detecting and dynamically resolving errors. Furthermore, Grotov et al. [45] developed an LLM-powered AI agent specifically designed to handle stateful and nonlinear notebook debugging cases, demonstrating that agent-based debugging can effectively resolve complex errors in computational notebooks. Their research also highlights the need to balance automation with human oversight, carefully manage interface complexity, and address security considerations (e.g., sandboxing).
@@ 5.5.3 Detecting Data Leakage
In data science projects, data leakage occurs when information from outside the training dataset is used to develop a model [155]. This issue can arise in Jupyter Notebook if code cells are executed in the wrong order. Although executing cells in an arbitrary sequence may seem harmless, it can lead to data leakage bugs, such as unintentionally sharing information between the training and test datasets. To provide warnings about potential data leakage, a static analysis framework has been developed that uses abstract interpretation for intracell static analysis, ensuring both efficiency and guaranteed termination [155]. Data leakage in ML can be detected early in Jupyter notebooks by turning dataframe operations across cells into a graph. Negrini et al. [103] showed that building such a graph-based model can capture dataframe operations across all notebook cells in their actual execution order, allowing them to detect data leakage effectively in Jupyter notebooks. Yang et al. [197] also tackled data leakage in data science notebooks by developing a static analysis tool that models how data flows across notebook cells. This tool tracks the relationships between datasets, transformations, and model evaluation steps to identify common leakage patterns specific to notebooks, such as the use of test data during preprocessing, the reuse of test data for model selection (known as multi-test leakage), and overlaps between training and test datasets.
Summary of testing and debugging Researchers have focused on developing automated tools designed explicitly for bug detection in notebooks. Those tools can deal with different types of bugs, for example, name-value inconsistencies [109], data leakage issues [197, 155], and output inconsistency [34]. Researchers explored LLM-powered solutions for debugging and resolving notebook errors [45, 47]. However, more research on notebook testing is needed to deal with other bugs such as performance and kernel-related bugs.
@@ 5.6 Visualization in Notebooks
As an important data analysis outcome, visualization is vital in Jupyter notebooks [145, 2]. It also serves software engineering purposes, as it can be used to validate the correctness of a notebook by allowing comparisons between expected and actual outcomes. Additionally, visualization is crucial for managing and enhancing communication within notebook projects.
@@ 5.6.1 Empirical Studies on Visualization in Notebooks
Visualization in notebooks can be divided into two main categories: data visualization and workflow visualization. Data visualization refers to the process of displaying the output of a code cell. One of the key features of Jupyter Notebook is the ability to visualize outputs alongside the code cells. Researchers showed that one out of four issues (1,071 out of 4,210) in the studied Jupyter Notebook projects contain at least one visualization-related issue [2]. The study also noticed that visual content is not limited to communicating user interaction design but contains various types of information, such as command-line content or code snippets. Settewong et al. [145] identified nine main reasons for visualizing data when notebook users write code. They analyzed 68 notebooks containing 821 visualizations and categorized them into nine types, showing that the most frequent use cases are for visualizing data distribution (e.g., histograms, box plots) and data frequency (e.g., bar charts, count plots) making these the core tools for understanding and communicating data insights during coding.
In addition, Wootton et al. [194] investigated the exploratory data analysis (EDA) process within Jupyter notebooks and identified distinct temporal and sequential patterns. They proposed quantitative metrics for measuring visualization usage, including revisit counts, representational diversity, and representation velocity. Wang et al. [184] showed that effective visual analytics in notebooks depends not only on what is visualized, but also on how visualizations are designed to fit the notebook environment. Their study identified four key design dimensions that shape the user experience, focusing on how visualizations connect with notebook code, where their data comes from, when they appear, and how easily they can be reused.
Another type of visualization in notebooks is workflow visualization, which helps users understand and navigate the structure of notebook code. Ramasamy et al. [127] introduced a tool called MARG, which visualizes data science notebooks as a graph of decisions, forks, and dead-ends instead of a linear list of cells. This approach reveals the nonlinear nature of real-world notebooks and allows users to trace the workflow more effectively. Their study showed that such visualizations significantly improve the understanding of the workflow of data science notebooks.
@@ 5.6.2 Interactive Visualization
Interactive visualizations in Jupyter Notebook refer to an interactive interface within notebooks that allows users to analyze data through interactive charts. These features greatly enhance the usability and practicality of Jupyter notebooks by facilitating real-time analysis, monitoring, and interaction with code and output [195, 77]. Interactive visualization also provides valuable insights and support informed decision-making by integrating visual analytics directly into the notebook environment. For instance, Scully-Allison et al. [144] developed a notebook-embedded interactive visualization that traces and synchronizes visual components with code cells. Furthermore, Guo et al. [50] developed bonXAI, which integrates interactive explainable AI (XAI) visualizations into Python-based ML workflows within notebooks.
Researchers have proposed interactive dashboards to help users explore and analyze their Jupyter notebooks. Kwon et al. [77] developed a customizable interactive dashboard to support data-centric NLP in Jupyter notebooks. Their system includes built-in text transformation operations and various visual analysis features, allowing users to create interactive dashboards seamlessly. Similarly, Wu et al. [195] introduced a flexible visualization dashboard that complements traditional code cells, offering users an intuitive interface for data interaction. By importing the necessary library, users can generate visualizations by clicking on dataset columns or creating custom data queries, such as scatter plots. This ensures real-time updates to code cells based on interactions with the dashboard.
Recent works have also explored extending notebook interaction beyond traditional desktop interfaces. In et al. [60] investigated adapting computational notebook interfaces into virtual reality (VR) and introduced embodied gestures for efficient non-linear exploration to enhance notebook navigation.
Summary of visualization in notebooks Visualization is essential in Jupyter Notebook for data analysis, result validation, and effective communication. Research shows that common applications include visualizing data distributions, statistical measures, and ML workflows [145, 2]. Researchers demonstrated that interactive visualizations enhance usability by enabling real-time interactions with both data and code [195, 77]. Several approaches have been proposed to trace and synchronize visual components with notebook code cells [144, 50]. Additionally, nonlinear workflow visualizations improve the comprehension of complex notebooks, and while established guidelines encourage integration, modularity to improve visual analytics [184, 127, 194].
@@ 5.7 Best Practices in Notebooks
It is essential to adhere to best practices in Jupyter notebooks to enhance code quality and collaborative efforts [149, 46]. Unlike conventional software development, Jupyter notebooks are often utilized for exploratory data analysis, often involving frequent adjustments and iterative code execution, resulting in deviations from traditional software development methodologies, such as waterfall or test-driven approaches [179]. This section underscores the importance of adhering to best practices in Jupyter notebooks, with a specific emphasis on ensuring code style consistency and addressing the unique challenges of collaborative usage, especially within the realm of ML projects.
@@ 5.7.1 Following Code Style Standards
One of the prevalent challenges in Jupyter notebooks is the inconsistency in code style arising from the absence of enforced coding standards. Researchers showed that notebooks have 1.4 times more stylistic issues than traditional Python scripts [46]. Another study experimentally demonstrated that Jupyter notebooks are inundated with poor-quality code, such as violations of recommended coding practices, unused variables, and deprecated functions [179]. Considering the knowledge-sharing nature of Jupyter notebooks, these poor coding practices might be propagated to future developers.
Researchers have studied the adherence to best practices for coding from various angles. Siddik and Bezemer [149] evaluated the adherence to the PEP-8 code style guidelines in ML code in Jupyter notebooks. Their findings indicate that ML notebooks generally exhibit lower code quality than non-ML notebooks, with notable discrepancies in how packages and libraries are managed. To address this, a static analysis tool called Pynblint [122, 123] was implemented to identify quality issues in Jupyter notebooks. In addition, Candela et al. [15] studied the quality of Jupyter Notebook projects published by GLAM (Galleries, Libraries, Archives, and Museums) institutions within the cultural heritage sector. Their evaluation criteria focused on documentation, code readability, and metadata usage, identifying that these areas need improvement for better traceability and overall quality.

### v2

@@ 5.5 Testing and Debugging
Jupyter notebooks are widely used for prototyping, offering rapid iteration and immediate feedback [fangohr2020testing]. However, testing in notebooks differs from traditional software testing, as it may require verifying code correctness at the cell level rather than treating the entire notebook as a single executable unit. Due to the lack of standardized testing practices, data scientists have adopted various approaches. For example, some embed test cases within the notebook, while others use separate test notebooks to verify the correctness [chattopadhyay2020whats]. This section describes studies related to the testing of notebooks, including empirical studies on bug analysis and bug detection.
@@ 5.5.1 Empirical Studies on Testing Notebooks
de2024bug empirically analyzed the problems and challenges of identifying, diagnosing and resolving bugs in Jupyter notebooks. They presented the first comprehensive study of bugs in Jupyter Notebook projects, analyzing bugs in GitHub repositories and Stack Overflow posts, and interviews with data scientists. The study identifies eight types of bugs that are found in notebooks along with their studies. shome2024understanding investigated feedback mechanisms within ML-focused Jupyter notebooks and presented a taxonomy of explicit (assertions) and implicit (print statements and last cell outputs) feedback approaches. Their study also provided practical recommendations to encourage explicit testing for minimizing technical debt and improving reproducibility. wang2024why analyzed crashes in ML notebooks and categorized predominant exceptions such as NameError and ValueError. In addition, they reported that some root causes of these exceptions are related to the features of Jupyter Notebook, such as out-of-order cell executions and errors propagating from previous cells.
@@
Processing bugs Issues in memory or computation, i.e., memory overflow, data loss, or performance degradation. [subotic2022static, negrini2023static, yang2022data, wang2024using, xin2022provenance]
Implementation bugs General coding errors that can cause incorrect results or runtime errors, i.e., syntax, logic, variables, or algorithms-related bugs. [patra2022nalin, brachmann2020your, grotov2024untangling, grotov2024debug]
Cell defect Bugs related to notebook cell renderings, i.e., code cells, Markdown, or output interactions. [fangohr2020testing]
Environments and settings bugs Problems from misconfigured environments or dependencies that interrupt notebook execution, i.e., missing libraries or lack of dependencies. [brachmann2020your]
@@ 5.5.2 Detecting Bugs
Detecting bugs or errors in Jupyter Notebook presents unique challenges due to nonlinear workflows and hidden states, distinguishing them from traditional programming environments [robinson2022error]. Notebook users rely heavily on manual trial-and-error debugging, rather than structured testing practices, with over 70% of cell re-executions driven by code fixing and cleaning [titov2025observing]. xin2022provenance developed a framework to detect anomalies and identify their root causes by combining provenance data with performance metrics. Since Jupyter notebooks do not offer a built-in debugger, the study also presents strategies for debugging Jupyter notebooks by identifying cell workflows and the root cause of errors. To make the bug detection easier, tang2025characterising investigated the underlying root causes of bugs in computational notebooks and identified three dominant sources: incorrect logical implementation, erroneous variable assignments, and file misconfigurations. Complementing these findings, jiang2025exploring further emphasized that configuration issues are among the most frequent root causes of notebook bugs, followed closely by improper API usage. Together, these studies highlight the fragile interplay between code logic, environment setup, and external dependencies in notebook-based workflows.
@@
Another bug-detecting study introduced a tool called Vizier, which helps debug notebooks by maintaining a complete version history of notebooks, cells, and datasets, while tracking potential errors through fine-grained provenance [brachmann2020your]. This tool maintains a complete version history for each notebook and allows reproducible data manipulation through a spreadsheet mode [brachmann2020your].
@@
Another method to detect potential bugs in notebooks is to focus on inconsistencies in variable assignments, output visualization, or logical operations within the notebook [patra2022nalin]. For example, output inconsistency occurs when the output of a notebook cell does not match the expected output. An implementation of such an inconsistency-based approach is the nbval notebook validation plugin [fangohr2020testing]. nbval detects bugs in Jupyter notebooks by automatically executing each code cell and comparing its output with previously stored results. If there is any deviation, such as an unexpected change in numerical values, formatting differences, or execution failures, nbval reports a test failure. wang2024using showed that traditional static analyzers frequently fail to detect runtime tensor shape mismatch bugs in ML codes (particularly TensorFlow) within notebooks. Their study demonstrates how incorporating runtime information improves static analysis capabilities for detecting such common ML-related notebook bugs.
@@
Another type of inconsistency bug is the name-value inconsistency, which occurs when the name of a variable does not accurately reflect its assigned value. Unlike traditional software development environments, notebooks allow for nonlinear execution and frequent variable reuse, increasing the risk of such name-value inconsistencies [patra2022nalin]. For instance, naming a variable log_file while storing a list of all files in a directory can cause a name-value inconsistency bug, as a developer might assume that log_file stores a single file name, but it actually contains a list. To avoid confusion, the variable could be renamed to log_files or log_file_list. patra2022nalin developed Nalin, an automated tool for detecting name-value inconsistencies. This tool uses a neural model to predict whether a variable’s name and its assigned value are consistent. Nalin employs dynamic program analysis alongside deep learning to monitor variable activities throughout the execution process.
@@
Recently, advancements in LLMs have created promising opportunities for automated error detection and debugging. grotov2024untangling performed a detailed analysis of prevalent notebook error patterns and proposed an iterative, LLM-based method for detecting and dynamically resolving errors. Furthermore, grotov2024debug developed an LLM-powered AI agent specifically designed to handle stateful and nonlinear notebook debugging cases, demonstrating that agent-based debugging can effectively resolve complex errors in computational notebooks. Their research also highlights the need to balance automation with human oversight, carefully manage interface complexity, and address security considerations (e.g., sandboxing).
@@ 5.5.3 Detecting Data Leakage
In data science projects, data leakage occurs when information from outside the training dataset is used to develop a model [subotic2022static]. This issue can arise in Jupyter Notebook if code cells are executed in the wrong order. Although executing cells in an arbitrary sequence may seem harmless, it can lead to data leakage bugs, such as unintentionally sharing information between the training and test datasets. To provide warnings about potential data leakage, a static analysis framework has been developed that uses abstract interpretation for intracell static analysis, ensuring both efficiency and guaranteed termination [subotic2022static]. Data leakage in ML can be detected early in Jupyter notebooks by turning dataframe operations across cells into a graph. negrini2023static showed that building such a graph-based model can capture dataframe operations across all notebook cells in their actual execution order, allowing them to detect data leakage effectively in Jupyter notebooks. yang2022data also tackled data leakage in data science notebooks by developing a static analysis tool that models how data flows across notebook cells. This tool tracks the relationships between datasets, transformations, and model evaluation steps to identify common leakage patterns specific to notebooks, such as the use of test data during preprocessing, the reuse of test data for model selection (known as multi-test leakage), and overlaps between training and test datasets.
Summary of testing and debugging Researchers have focused on developing automated tools designed explicitly for bug detection in notebooks. Those tools can deal with different types of bugs, for example, name-value inconsistencies [patra2022nalin], data leakage issues [yang2022data, subotic2022static], and output inconsistency [fangohr2020testing]. Researchers explored LLM-powered solutions for debugging and resolving notebook errors [grotov2024debug, grotov2024untangling]. However, more research on notebook testing is needed to deal with other bugs such as performance and kernel-related bugs.
@@ 5.6 Visualization in Notebooks
As an important data analysis outcome, visualization is vital in Jupyter notebooks [settewong2022visualize, agrawal2022understanding]. It also serves software engineering purposes, as it can be used to validate the correctness of a notebook by allowing comparisons between expected and actual outcomes. Additionally, visualization is crucial for managing and enhancing communication within notebook projects.
@@ 5.6.1 Empirical Studies on Visualization in Notebooks
Visualization in notebooks can be divided into two main categories: data visualization and workflow visualization. Data visualization refers to the process of displaying the output of a code cell. One of the key features of Jupyter Notebook is the ability to visualize outputs alongside the code cells. Researchers showed that one out of four issues (1,071 out of 4,210) in the studied Jupyter Notebook projects contain at least one visualization-related issue [agrawal2022understanding]. The study also noticed that visual content is not limited to communicating user interaction design but contains various types of information, such as command-line content or code snippets. settewong2022visualize identified nine main reasons for visualizing data when notebook users write code. They analyzed 68 notebooks containing 821 visualizations and categorized them into nine types, showing that the most frequent use cases are for visualizing data distribution (e.g., histograms, box plots) and data frequency (e.g., bar charts, count plots) making these the core tools for understanding and communicating data insights during coding.
In addition, wootton2024charting investigated the exploratory data analysis (EDA) process within Jupyter notebooks and identified distinct temporal and sequential patterns. They proposed quantitative metrics for measuring visualization usage, including revisit counts, representational diversity, and representation velocity. wang2023supernova showed that effective visual analytics in notebooks depends not only on what is visualized, but also on how visualizations are designed to fit the notebook environment. Their study identified four key design dimensions that shape the user experience, focusing on how visualizations connect with notebook code, where their data comes from, when they appear, and how easily they can be reused.
Another type of visualization in notebooks is workflow visualization, which helps users understand and navigate the structure of notebook code. ramasamy2023visualising introduced a tool called MARG, which visualizes data science notebooks as a graph of decisions, forks, and dead-ends instead of a linear list of cells. This approach reveals the nonlinear nature of real-world notebooks and allows users to trace the workflow more effectively. Their study showed that such visualizations significantly improve the understanding of the workflow of data science notebooks. Expanding on this, NoteFlow by tian2025noteflow introduced a chart recommendation system for tracing data flows across notebook cells. NoteFlow can parse data flow, recommends context-aware charts, and enables forward/backward tracing of visual states, helping users monitor evolving data tables and identify anomalies during exploratory data analysis. Additionally wenskovitch2019albireo offered another graph visualization of notebook structure, encoding markdown and code cells as nodes and highlighting variable dependencies.
@@ 5.6.2 Interactive Visualization
Interactive visualizations in Jupyter Notebook refer to an interactive interface within notebooks that allows users to analyze data through interactive charts. These features greatly enhance the usability and practicality of Jupyter notebooks by facilitating real-time analysis, monitoring, and interaction with code and output [wu2020b2, kwon2023weedle]. Interactive visualization also provides valuable insights and support informed decision-making by integrating visual analytics directly into the notebook environment. For instance, scully2024design developed a notebook-embedded interactive visualization that traces and synchronizes visual components with code cells. Furthermore, guo2024explainability developed bonXAI, which integrates interactive explainable AI (XAI) visualizations into Python-based ML workflows within notebooks. ono2021interactive presented three core approaches for embedding interactivity, named as, matplotlib callbacks, visualization toolkits like Plotly and Altair, and custom HTML/JavaScript visualizations, which enable users to tailor visualizations to their analytical needs.
Researchers have proposed interactive dashboards to help users explore and analyze their Jupyter notebooks. kwon2023weedle developed a customizable interactive dashboard to support data-centric NLP in Jupyter notebooks. Their system includes built-in text transformation operations and various visual analysis features, allowing users to create interactive dashboards seamlessly. Similarly, wu2020b2 introduced a flexible visualization dashboard that complements traditional code cells, offering users an intuitive interface for data interaction. By importing the necessary library, users can generate visualizations by clicking on dataset columns or creating custom data queries, such as scatter plots. cai2025jupyter developed a toolkit for visualizing notebooks’ activity, offering real-time dashboards that show cell-level engagement, execution errors, and time-on-task metrics. These ensure real-time updates to code cells based on interactions with the dashboard.
Recent works have also explored extending notebook interaction beyond traditional desktop interfaces. in2024evaluating investigated adapting computational notebook interfaces into virtual reality (VR) and introduced embodied gestures for efficient non-linear exploration to enhance notebook navigation.
Summary of visualization in notebooks Visualization is essential in Jupyter Notebook for data analysis, result validation, and effective communication. Research shows that common applications include visualizing data distributions, statistical measures, and ML workflows [settewong2022visualize, agrawal2022understanding, tian2025noteflow]. Researchers demonstrated that interactive visualizations enhance usability by enabling real-time interactions with both data and code [wu2020b2, kwon2023weedle]. Several approaches have been proposed to trace and synchronize visual components with notebook code cells [scully2024design, guo2024explainability, ono2021interactive]. Additionally, nonlinear workflow visualizations improve the comprehension of complex notebooks, and while established guidelines encourage integration, modularity to improve visual analytics [wang2023supernova, ramasamy2023visualising, wootton2024charting].
@@ 5.7 Best Practices in Notebooks
It is essential to adhere to best practices in Jupyter notebooks to enhance code quality and collaborative efforts siddik2023code, grotov2022large. Unlike conventional software development, Jupyter notebooks are often utilized for exploratory data analysis, often involving frequent adjustments and iterative code execution, resulting in deviations from traditional software development methodologies, such as waterfall or test-driven approaches wang2020better. This section underscores the importance of adhering to best practices in Jupyter notebooks, with a specific emphasis on ensuring code style consistency and addressing the unique challenges of collaborative usage, especially within the realm of ML projects.
@@ 5.7.1 Following Code Style Standards
One of the prevalent challenges in Jupyter notebooks is the inconsistency in code style arising from the absence of enforced coding standards. Researchers showed that notebooks have 1.4 times more stylistic issues than traditional Python scripts grotov2022large. Research found that Jupyter notebooks are generally larger, but less complex, and tend to use built-in functions much more than scripts and have a higher coupling rate [adams2023comparison]. Another study experimentally demonstrated that Jupyter notebooks are inundated with poor-quality code, such as violations of recommended coding practices, unused variables, and deprecated functions wang2020better. Considering the knowledge-sharing nature of Jupyter notebooks, these poor coding practices might be propagated to future developers.
Lightweight software engineering practices are adopted in Jupyter notebooks to improve clarity, reproducibility, and long-term reuse of research code [huang2025scientists]. Such practices help balance the need for exploration with the need for organized and understandable code. In their study, huang2025scientists identified eight key quality goals, such as clarity, reproducibility, and reusability, and suggested 18 practical tactics to achieve them. Some common tactics include keeping one main task per cell, using markdown cells to explain code and results, and improving or cleaning up code when necessary. Similarly, cai2025jupyterNotebook found that notebooks with well-defined and consistently used variables and functions make it easier for others to understand the workflow. This structured approach helps developers and researchers work together more effectively and maintain a smooth flow of collaboration.
Prior work has examined adherence to coding best practices from multiple perspectives. For example, siddik2023code evaluated the adherence to the PEP-8 code style guidelines in ML code in Jupyter notebooks. Their findings indicate that ML notebooks generally exhibit lower code quality than non-ML notebooks, with notable discrepancies in how packages and libraries are managed. To address this, a static analysis tool called Pynblint [quaranta2022pynblint, quaranta2024pynblint] was implemented to identify quality issues in Jupyter notebooks. In addition, candela2023approach studied the quality of Jupyter Notebook projects published by GLAM (Galleries, Libraries, Archives, and Museums) institutions within the cultural heritage sector. Their evaluation criteria focused on documentation, code readability, and metadata usage, identifying that these areas need improvement for better traceability and overall quality.